OpenStudy (anonymous):

You have 80 yards of fencing to enclose a rectangular region. Find the dimensions of the rectangle that maximize the enclosed are. What is the maximum area?

OpenStudy (saifoo.khan):

400

OpenStudy (anonymous):

how did u get it..

OpenStudy (saifoo.khan):

i have did that before. i know the answer. :p

OpenStudy (saifoo.khan):

the method is sooo long. :9

OpenStudy (saifoo.khan):

:(

OpenStudy (anonymous):

lol wow.. man i wanna see the steps. can u try atleast

OpenStudy (anonymous):

Yeah it is long.. Just try.

OpenStudy (saifoo.khan):

:/

OpenStudy (saifoo.khan):

it's mathlab right?

OpenStudy (saifoo.khan):

let me write the method.

OpenStudy (saifoo.khan):

Square is the rectangular shape that maximizes area. perimeter = 4L 80 /4 = L L = 20 Area = 20^2 = 400

OpenStudy (saifoo.khan):

i did a shortcut.

OpenStudy (anonymous):

this is what am asked to find. A.) Find vertex B.) Find dimesion C.) Find maximum area

OpenStudy (anonymous):

which is which?

OpenStudy (saifoo.khan):

u asked for max area.. it is 400.

OpenStudy (saifoo.khan):

now u need vertex and dimension?

OpenStudy (anonymous):

how do i find them?

OpenStudy (saifoo.khan):

finding..

OpenStudy (anonymous):

okay

OpenStudy (saifoo.khan):

WAIT! which week is that?

OpenStudy (anonymous):

week?

OpenStudy (saifoo.khan):

im stuck with the vertex! :O

OpenStudy (anonymous):

ch.3.1 if thats wat u mean by week?

OpenStudy (saifoo.khan):

no, i mean which checkpoint or quiz..

OpenStudy (anonymous):

3.1 apllication exercises..

OpenStudy (anonymous):

its like the review part of section 3.1

OpenStudy (anonymous):

polynomial and rational functions

OpenStudy (saifoo.khan):

let me find it..

OpenStudy (saifoo.khan):

aren't u doing mathlab? do u?

OpenStudy (anonymous):

college algebra

OpenStudy (saifoo.khan):

i know that's College Algebra!! but i can't figure it out in which week that question is asked!

OpenStudy (anonymous):

am lost with the week part. all i can say am in exercise 3.1 question # 68 chapter 3 polynomial and rational functions.

OpenStudy (saifoo.khan):

im sorry, but i can't figure out the the vertex part. :/

OpenStudy (anonymous):

:( can u give an example of how u would use week? so i can see if i know

OpenStudy (saifoo.khan):

are u in UoP or something or are u doing mathlab?

OpenStudy (anonymous):

UoP?? n nah no mathlab

OpenStudy (saifoo.khan):

Univ of Pheonix.

OpenStudy (saifoo.khan):

what if u skip this and move to others?

OpenStudy (anonymous):

okay did figure out the dimensions

OpenStudy (saifoo.khan):

Nope, the vertex will give us that. :(

OpenStudy (anonymous):

like all the questions i have are asking for vertex

OpenStudy (saifoo.khan):

what are others?

OpenStudy (anonymous):

You have 200 feet of fencing to enclose a rectangular plot that borders on a river. If you do not fence the side along the river, find the length and width of the plot that will maximize the area. What is the largest are that can be enclosed? A.) Find Vertex B.) Find dimensions L and W C.) Find maximum area

OpenStudy (anonymous):

Among all pairs of numbers whose difference is 24, find a pair whose product is as small possible. What is the minimum product? A.) Find Vertex B.) Find numbers C.) Find minimum product

OpenStudy (saifoo.khan):

is that calculus? :O

OpenStudy (anonymous):

nope college algerbra: an early functions approach

OpenStudy (saifoo.khan):

Wait, i can solve that by looking at the example! :D im looking on page 307

OpenStudy (anonymous):

Yess!!!

OpenStudy (saifoo.khan):

P = xy x - y = 24

OpenStudy (anonymous):

Solvinng??

OpenStudy (saifoo.khan):

why not.

OpenStudy (anonymous):

Are you working on it?

OpenStudy (saifoo.khan):

Vertex = (24,0)

OpenStudy (anonymous):

how bout B?

OpenStudy (saifoo.khan):

im doing this one: Among all pairs of numbers whose difference is 24, find a pair whose product is as small possible. What is the minimum product? A.) Find Vertex B.) Find numbers C.) Find minimum product A) (24,0) B) i think it's 1 and -24.

OpenStudy (anonymous):

C.

OpenStudy (saifoo.khan):

trying to solve..

OpenStudy (saifoo.khan):

If u can wait for a day, i can go through all these and solve your questions tommorw! how much time do u have? these are easy, i just need some time to go through all this!

OpenStudy (anonymous):

ohh man only today. class tomorrow.. it a total of 6 question but 2 of each are the same. so i u can answer 3 of them i can do tha other 3

OpenStudy (saifoo.khan):

ok, let me read again then..

OpenStudy (saifoo.khan):

x^2 -24x a = 1 b = -24

OpenStudy (saifoo.khan):

function will have a minimum point. -b/2a -(-24)/2(1) 12 This means that the product, P, of two numbers whose difference is 24 is a minimum when one of the numbers, x, is 12 .

OpenStudy (saifoo.khan):

y = x - 24 y = 12-24 y = -12

OpenStudy (saifoo.khan):

The number pair whose difference is 24 and whose product is as small as possible is 12, -12. The minimum product is 12(-12) , or -144

OpenStudy (saifoo.khan):

Correct answer! :D

OpenStudy (anonymous):

Wow that was fast

OpenStudy (saifoo.khan):

Thank you! :D

OpenStudy (saifoo.khan):

next?

OpenStudy (saifoo.khan):

post that on left please, this is pretty long thread.

OpenStudy (anonymous):

okay so did u answer al A,B and C and how do i post it on the left?

OpenStudy (saifoo.khan):

Yes, that's the complete solution! :D

OpenStudy (saifoo.khan):

i mean on the left pane.. where other questions are asked.